HERA Beam Dynamics Meeting / Minutes : 09.10.2003
D.Barber, M.Bieler, W.Decking, S.Herb,
M.Hoffmann, J.Keil, M.Minty, D.Pitzl, M.Vogt, R.Wanzenberg,
- Venting of flange SR11 : orbit controll
- apearently horizontal offset of +20mm in GI/GJ SR !!
- immediate preventive actions :
- PT100's installed at downstream flanges of absorber 4 (SR11)
- console program :
submenue of local-history-viwer (M.Lomperski)
- short-term plans :
- new, more stringent procedures for assigning files
- tracking/ energy disposition simulations
- vacuum seems to recover faster than expected
- Hot spot OR34 :
- last weekend : hot spot (35μS) of 5-10cm length found on plain
beam pipe OR34. [20cm left of bellow and 30cm right of getter pump]
- aperture in that region seems to be reduced
- however, injection OK, lifteime OK
- no obvious obstacle found
- later today the aperture will be measured with beam
- QC's OR12/OR14 :
The QC's have been lowered. Improvement not yet veryfied with beam.
⇒ later today
- Vakuum leak at Cavities OR 75/76 :
- all 4 plungers changed
- since then : vacuum conditioning with RF on and warercooling off
- water leak found in one of the old plungers
- 04:00 this morning : RF off and ``adiabtic'' cool down ;
then water on
- next : condition cavity vaccum during operation
- Short to ground at BU NL :
- short still not better localizes than BU NL section (most likely
center magnet or one towards hall N)
- PS now operated with virtual ground moved from symmetry point (O)
to BU NL
- ramp & fast dump tests finished. Result :
temporary operation in this mode up to 920GeV seems OK
- mid-/long-term plans : buy some spare coils and start replacing
the most suspicious looking and obviously defect coils
--- until then : let's hope for the best ! ---
- NMR / Zanon reference magnet :
During thies weeks p-Injektion studies, the magnet compensation
(NMR inside reference magnets hall W) seemed to be unstable.
Experts are working on the problem.
- Cable length compensation for HIT to be switched to ``winter''
eitehr this or next week
- Agenda for next week's meting
orbit analysis / orbit steering / orbit interlock /
file assignment procedures / ...
HERA BeamDynamicsMeeting homepage,
DESY homepage ,
MPY homepage ,
Machine Status .